OpenStudy (anonymous):

I forgot how to do this. -(3x - 2) = -4(x + 1) - 4

pooja195 (pooja195):

-(3x - 2) = -4(x + 1) - 4 Start by distributing the negative

OpenStudy (anonymous):

-3x +2 = -4x -4 -4 -3x +2 = -4x -8 2 = -x - 8 10 = -x x = -10?

pooja195 (pooja195):

Correct well done :)

OpenStudy (anonymous):

Thanks! One more? I have an idea of what it is. x = 32? I multiply it right?

pooja195 (pooja195):

Correct

OpenStudy (anonymous):

Alright, thanks!

pooja195 (pooja195):

No problem let me know if you need more help :)

OpenStudy (anonymous):

I'll try to do every problem on my own first and post my answer, and I'll look for judgement. :) I can give you more medals that way.

pooja195 (pooja195):

I dont care about medals I just want you to learn! :)